# Multi-color characterization of optically invisible FU Orionis-type outbursts: Demonstration and prospects for the WINTER survey

### v1

@@ 3.4 Luminosities and mass accretion rates
with as the adopted inner disk truncation radius (Gullbring et al., 1998). Applying this relation, we find that in WNTR24-cua, measured approximately 7 years after the onset of its outburst, corresponds to an accretion rate of . For WNTR24-egv, the larger increase of , derived from observations taken roughly 10 years after the outburst began, implies .
@@
Given the uncertainty in the distance to WNTR24-cua, which ranges from kpc (calculated in Section 3.1) to the two proposed values of and kpc from Guo et al. (2024a), we scale the luminosities and accretion rates accordingly. At 3.0 kpc, the post-outburst luminosity increases to and the inferred accretion rate to . At 6.1 kpc, these values become and . These values remain within the expected range for FUors but suggest that WNTR24-cua may be more luminous than implied under the nominal 2.4 kpc assumption.
4 Discussion and Summary
WINTER identified two new FUor outbursts that are bright in the near-infrared, also detected in the mid-infrared, but largely obscured at optical wavelengths. WNTR24-cua is a previously characterized FUor and WNTR24-egv is a newly classified FUor from this sample. Long-term monitoring with surveys like WISE and Gaia show the rise and color evolution of the bursts, revealing years-long events with slow declines. WINTER observations of the events confirmed that these objects were bright in the NIR and prompted rapid spectroscopic follow-up. The spectra for both events show cool objects with mixed spectral types with evidence of winds or outflows. Furthermore, the spectra confirm the events as FUors in the Connelley & Reipurth 2018 classification scheme. Compared to the larger sample of FUors, the two WINTER events show notably high CO equivalent widths (Figure 4) and steep spectral slopes.
@@ 4 Discussion and Summary
Overall, these studies highlight the importance of multi-color monitoring of the Galactic plane to understand the triggering mechanisms of FUors. The classical thermal instability model predicts that the IR rise lags behind the optical (Bell et al., 1995). In contrast, magnetorotational instability (MRI) models predict an infrared precursor, with the delay depending on the luminosity and mass accretion rate. For the largest outbursts—such as FU Orionis itself—Cleaver et al. (2023) argue that the outburst originates far out in the disk, producing an infrared precursor decades before the optical rise. However, the delay is shorter for smaller peak outbursts, where the instability occurs closer to the star. Models by Nayakshin & Elbakyan (2024) instead suggest that the evaporation of a close-in planet can trigger FUor outbursts; in this case, the delay does not depend on the mass accretion rate. Guo et al. (2024b) find that such a scenario is more consistent with similar delays observed in outbursts with varying peak accretion rates.
@@
5 Acknowledgments
WINTER’s construction is made possible by the National Science Foundation under MRI grant number AST-1828470 with early operations supported by AST-1828470. Significant support for WINTER also comes from the California Institute of Technology, the Caltech Optical Observatories, the Bruno Rossi Fund of the MIT Kavli Institute for Astrophysics and Space Research, the David and Lucille Packard Foundation, and the MIT Department of Physics and School of Science. D.F.’s contribution to this material is based upon work supported by the National Science Foundation under Award No. AST-2401779. This research award is partially funded by a generous gift of Charles Simonyi to the NSF Division of Astronomical Sciences. The award is made in recognition of significant contributions to Rubin Observatory’s Legacy Survey of Space and Time. Furthermore, we acknowledge the support of the National Aeronautics and Space Administration through ADAP grant number 80NSSC24K0663. This work is based on observations obtained at the Infrared Telescope Facility, which is operated by the University of Hawai‘i under contract 80HQTR19D0030 with the National Aeronautics and Space Administration. We acknowledge ESA Gaia, DPAC and the Photometric Science Alerts Team (http://gsaweb.ast.cam.ac.uk/alerts). This research made use of hips2fits,333https://alasky.cds.unistra.fr/hips-image-services/hips2fits a service provided by CDS.

### v2

@@ III.4 Luminosities and mass accretion rates
with as the adopted inner disk truncation radius (Gullbring et al., 1998). Applying this relation, we find that in WNTR24-cua, measured approximately 7 years after the onset of its outburst, corresponds to an accretion rate of . For WNTR24-egv, the larger increase of , derived from observations taken roughly 10 years after the outburst began, implies . These derived values are summarized in Table 1.
@@
Given the uncertainty in the distance to WNTR24-cua, which ranges from kpc (calculated in Section III.1) to the two proposed values of and kpc from Guo et al. (2024a), we scale the luminosities and accretion rates accordingly. At 3.0 kpc, the post-outburst luminosity increases to and the inferred accretion rate to . At 6.1 kpc, these values become and . These values remain within the expected range for FUors but suggest that WNTR24-cua may be more luminous than implied under the nominal 2.4 kpc assumption.
IV Discussion and Summary
WINTER identified two new FUor outbursts that are bright in the near-infrared, also detected in the mid-infrared, but largely obscured at optical wavelengths. WNTR24-cua is a previously characterized FUor and WNTR24-egv is a newly classified FUor from this sample. Long-term monitoring with surveys like WISE and Gaia show the rise and color evolution of the bursts, revealing years-long events with slow declines. WINTER observations of the events confirmed that these objects were bright in the NIR and prompted rapid spectroscopic follow-up. The spectra for both events show cool objects with mixed spectral types with evidence of winds or outflows. Furthermore, the spectra confirm the events as FUors in the Connelley and Reipurth 2018 classification scheme. Compared to the larger sample of FUors, the two WINTER events show notably high CO equivalent widths (Figure 4) and steep spectral slopes.
@@ IV Discussion and Summary
Overall, these studies highlight the importance of multi-color monitoring of the Galactic plane to understand the triggering mechanisms of FUors. The classical thermal instability model predicts that the IR rise lags behind the optical (Bell et al., 1995). In contrast, magnetorotational instability (MRI) models predict an infrared precursor, with the delay depending on the luminosity and mass accretion rate. For the largest outbursts—such as FU Orionis itself—Cleaver et al. (2023) argue that the outburst originates far out in the disk, producing an infrared precursor decades before the optical rise. However, the delay is shorter for smaller peak outbursts, where the instability occurs closer to the star. Models by Nayakshin and Elbakyan (2024) instead suggest that the evaporation of a close-in planet can trigger FUor outbursts; in this case, the delay does not depend on the mass accretion rate. Guo et al. (2024b) find that such a scenario is more consistent with similar delays observed in outbursts with varying peak accretion rates.
@@
V Acknowledgments
WINTER’s construction is made possible by the National Science Foundation under MRI grant number AST-1828470 with early operations supported by AST-1828470. Significant support for WINTER also comes from the California Institute of Technology, the Caltech Optical Observatories, the Bruno Rossi Fund of the MIT Kavli Institute for Astrophysics and Space Research, the David and Lucille Packard Foundation, and the MIT Department of Physics and School of Science. D.F.’s contribution to this material is based upon work supported by the National Science Foundation under Award No. AST-2401779. This research award is partially funded by a generous gift of Charles Simonyi to the NSF Division of Astronomical Sciences. The award is made in recognition of significant contributions to Rubin Observatory’s Legacy Survey of Space and Time. Furthermore, we acknowledge the support of the National Aeronautics and Space Administration through ADAP grant number 80NSSC24K0663. This work is based on observations obtained at the Infrared Telescope Facility, which is operated by the University of Hawai‘i under contract 80HQTR19D0030 with the National Aeronautics and Space Administration. We acknowledge ESA Gaia, DPAC and the Photometric Science Alerts Team (http://gsaweb.ast.cam.ac.uk/alerts). This research made use of hips2fits,333https://alasky.cds.unistra.fr/hips-image-services/hips2fits a service provided by CDS. Any use of generative AI in this manuscript adheres to ethical guidelines for use and acknowledgment of generative AI in academic research. Each author has made a substantial contribution to the work, which has been thoroughly vetted for accuracy, and assumes responsibility for the integrity of their contributions.